OpenStudy (anonymous):

Tom and Jerry have a mowing business. It takes Tom 30minutes to mow a yard and it takes Jerry 40minutes to mow the same yard. If they worked together, how long will it take them to mow this yard? (Average Rates of Change topic)

OpenStudy (lgbasallote):

hintl \[\huge \frac t{30} + \frac t{40} = 1\] solve for t

OpenStudy (lgbasallote):

that would be the answer

OpenStudy (anonymous):

How did you get that equation?

OpenStudy (anonymous):

I got the answer 120/7 which is correct but how did you come up with that equation? Why does it equal 1?lgbasallote

OpenStudy (phi):

I think of these problems as rate*time = job (almost like rate*time=distance) rate is measured in job per amount of time so Tom works at a rate of 1 job/30 mins Jerry at a rate of 1 job per 40 mins or 1/40 job/mins together the rate is (1/30 + 1/40) and the equation is (1/30 + 1/40)*t = 1 job solve for t (time in minutes)

OpenStudy (anonymous):

Wow, that's neat. Thanks :) That makes much more sense.